# Deformation of Fuchsian representations and proper affine actions

### v1

@@ 5.3. Deformation of surface groups
In this subsection we assume that $\Gamma$ is the fundamental group of a compact surface without boundary. Let $\iota:\mathsf{SL}_{2}\to\mathsf{SL}_{n}$ be the irreducible representation. Then any injective representation $\rho:\Gamma\to\mathsf{SL}_{n}$ is called a Fuchsian representation if and only if $\rho=\iota\circ\tau$ for some $\tau:\Gamma\to\mathsf{SL}_{2}$. Moreover, any representation $\rho:\Gamma\to\mathsf{SL}_{n}$ which can be continuously deformed to obtain a Fuchsian representation is called a Hitchin representation. We denote the space of Hitchin representations by $\mathsf{Hit}(\Gamma,\mathsf{SL}_{n})$. Suppose $\rho$ be a Fuchsian representation and $\mathsf{T}_{[\rho]}\mathsf{Hit}(\Gamma,\mathsf{SL}_{n})$ be the tangent space at $[\rho]$.
@@
Suppose $\{\rho_{t}\mid\rho_{t}:\Gamma\to\mathsf{SL}_{n},t\in(-1,1)\}$ is an analytic one parameter family of Hitchin representations whose tangent at $\rho=\rho_{0}$ is given by the cocycle $u:\Gamma\to\mathfrak{sl}_{n}$. We use results from Section 6.3 of Bridgeman–Canary–Labourie–Sambarino [BCLS15] to obtain an analytic family of functions $f_{t}:\mathsf{U}\Gamma\to\mathfrak{sl}_{n}^{0}$ such that
$\int_{\gamma}f_{t}=\frac{\mathtt{Jd}_{\rho_{t}}(\gamma)}{\ell(\gamma)}.$
Hence, $\alpha(f_{t})$ also varies analytically. It follows that for some $g_{t}:\mathsf{U}\Gamma\to\mathfrak{sl}_{n}^{0}$,
$\alpha(f_{t})=\alpha(f_{0})+t\alpha(f^{\prime}_{0})+\frac{t^{2}}{2!}\alpha(g_{t}).$
Now using Proposition 4.2.4 we obtain that for all $\gamma\in\Gamma$ the following holds:
$\int_{\gamma}\alpha(f^{\prime}_{0})=\alpha\left(\frac{\mathtt{M}_{\rho,u}(\gamma)}{\ell(\gamma)}\right)\leq-c<0.$
We use Lemmas 3 and 6 of Goldman–Labourie [GL12] to observe that we can choose $f_{t}$ such that $\alpha(f^{\prime}_{0})<0$. As $\mathsf{U}\Gamma$ is compact, there exists some $c_{1}>0$ and $b_{t}>0$ such that $\alpha(f^{\prime}_{0})\leq-c_{1}$ and $\alpha(g_{t})\leq b_{t}$. Therefore, for small enough $t>0$, there exists $c_{2}>0$ such that
$\alpha(f_{t})\leq\alpha(f_{0})-tc_{2}\leq\alpha(f_{0}).$
Hence, for all $\gamma\in\Gamma$ we have $\alpha(\mathtt{Jd}_{\rho_{t}}(\gamma))\leq\alpha(\mathtt{Jd}_{\rho}(\gamma))$ and we deduce that
$\lim_{s\to\infty}\frac{1}{s}\log|\{\gamma\mid\alpha(\mathtt{Jd}_{\rho}(\gamma))\leq s\}|\leq\lim_{s\to\infty}\frac{1}{s}\log|\{\gamma\mid\alpha(\mathtt{Jd}_{\rho_{t}}(\gamma))\leq s\}|.$
Now using Corollary 1.4 of Potrie–Sambarino [PS17] we deduce that $\rho_{t}$ is a Fuchsian representation for all small enough $t>0$. Finally, we use results from Goldman–Margulis [GM00] to obtain that $0\in\mathtt{M}\text{-}\mathsf{Spec}(\rho,u)$. Hence, $0\in\alpha\left(\mathtt{M}\text{-}\mathsf{Spec}(\rho,u)\right)$, a contradiction to our initial assumption. Therefore, our result follows. ∎

### v3

@@ 5.3. Deformation of surface groups
In this subsection we assume that $\Gamma$ is the fundamental group of a compact surface without boundary of genus atleast two. Let $\iota:\mathsf{SL}_{2}\to\mathsf{SL}_{n}$ be the irreducible representation. Then any injective representation $\rho:\Gamma\to\mathsf{SL}_{n}$ is called a Fuchsian representation if and only if $\rho=\iota\circ\tau$ for some $\tau:\Gamma\to\mathsf{SL}_{2}$. Moreover, any representation $\rho:\Gamma\to\mathsf{SL}_{n}$ which can be continuously deformed to obtain a Fuchsian representation is called a Hitchin representation. We denote the space of Hitchin representations by $\mathsf{Hit}(\Gamma,\mathsf{SL}_{n})$. Suppose $\rho$ be a Fuchsian representation and $\mathsf{T}_{[\rho]}\mathsf{Hit}(\Gamma,\mathsf{SL}_{n})$ be the tangent space at $[\rho]$.
@@
Suppose $\{\rho_{s}\mid\rho_{s}:\Gamma\to\mathsf{SL}_{n},s\in(-1,1)\}$ is an analytic one parameter family of Hitchin representations whose tangent at $\rho=\rho_{0}$ is given by the cocycle $u:\Gamma\to\mathfrak{sl}_{n}$. We use results from Section 6.3 of Bridgeman–Canary–Labourie–Sambarino [BCLS15] to obtain an analytic family of functions $f_{s}:\mathsf{U}\Gamma\to\mathfrak{sl}_{n}^{0}$ such that
$\int_{\gamma}f_{s}=\frac{\mathtt{Jd}_{\rho_{t}}(\gamma)}{t_{\gamma}}.$
Hence, $\alpha(f_{s})$ also varies analytically. It follows that for some $g_{s}:\mathsf{U}\Gamma\to\mathfrak{sl}_{n}^{0}$,
$\alpha(f_{s})=\alpha(f_{0})+s\alpha(f^{\prime}_{0})+\frac{s^{2}}{2!}\alpha(g_{s}).$
Now using Proposition 4.2.4 and Corollary 8.3H of Gromov [Gro87] we obtain some $c_{0}>0$ such that for all $\gamma\in\Gamma$ the following holds:
$\int_{\gamma}\alpha(f^{\prime}_{0})=\alpha\left(\frac{\mathtt{M}_{\rho,u}(\gamma)}{t_{\gamma}}\right)\leq-c_{0}<0.$
We use Lemmas 3 and 6 of Goldman–Labourie [GL12] to observe that we can choose $f_{s}$ such that $\alpha(f^{\prime}_{0})<0$. As $\mathsf{U}\Gamma$ is compact, there exists some $c_{1}>0$ and $b_{s}>0$ such that $\alpha(f^{\prime}_{0})\leq-c_{1}$ and $\alpha(g_{s})\leq b_{s}$. Therefore, for small enough $s>0$, there exists $c_{2}>0$ such that
$\alpha(f_{s})\leq\alpha(f_{0})-sc_{2}\leq\alpha(f_{0}).$
Hence, for all $\gamma\in\Gamma$ we have $\alpha(\mathtt{Jd}_{\rho_{s}}(\gamma))\leq\alpha(\mathtt{Jd}_{\rho}(\gamma))$ and we deduce that
$\lim_{t\to\infty}\frac{1}{t}\log|\{\gamma\mid\alpha(\mathtt{Jd}_{\rho}(\gamma))\leq t\}|\leq\lim_{t\to\infty}\frac{1}{t}\log|\{\gamma\mid\alpha(\mathtt{Jd}_{\rho_{s}}(\gamma))\leq t\}|.$
Now using Corollary 1.4 of Potrie–Sambarino [PS17] we deduce that $\rho_{s}$ is a Fuchsian representation for all small enough $s>0$. Finally, we use results from Goldman–Margulis [GM00] to obtain that $0\in\mathtt{M}\text{-}\mathsf{Spec}(\rho,u)$. Hence, $0\in\alpha\left(\mathtt{M}\text{-}\mathsf{Spec}(\rho,u)\right)$, a contradiction to our initial assumption. Therefore, our result follows. ∎
@@
Remark 5.3.7.
The above proposition can also be proved more directly by considering the complete reducibility of $\mathfrak{sl}_{2}$ representations and Theorem 1.1 of Labourie [Lab01]. We thank the anonymous referee for bringing this to our notice.
Notation 5.3.8.
We use the following notation:
$\mathsf{T}_{odd}:=\bigoplus_{m=1}^{\lfloor\frac{n-1}{2}\rfloor}\mathsf{T}_{2m+1}\text{ and }\mathsf{T}_{even}:=\bigoplus_{m=1}^{\lfloor\frac{n}{2}\rfloor}\mathsf{T}_{2m}.$
Proposition 5.3.9.
Suppose $(\rho,u):\Gamma\to\mathsf{SL}_{n}\ltimes\mathfrak{sl}_{n}$ is an injective homomorphism such that $\rho$ is a Fuchsian representation and $u\in\mathsf{T}_{odd}$. Then $(\rho,u)(\Gamma)$ does not act properly on $\mathfrak{sl}_{n}$.
Proof.
As $u$ corresponds to an odd differential, using Equation (5), Propositions 2.1.1 and 4.0.2 of Labourie–Wentworth [LW18] we obtain that
$\mathtt{M}_{\rho,u}\left(\gamma^{-1}\right)=-\mathtt{M}_{\rho,u}(\gamma)$
for all $\gamma\in\Gamma$. Hence, for any $\gamma,\eta\in\Gamma$ we have
$\mathtt{M}_{\rho,u}\left(\eta\gamma^{n}\eta^{-1}\right)+\mathtt{M}_{\rho,u}\left(\gamma^{-n}\right)=0.$
We choose the elements $\gamma$ and $\eta$ from $\Gamma$ such that they are of infinite order and transverse. It follows that, we have $\{\eta\gamma^{n}\eta^{-1}\gamma^{-n}\}_{n=0}^{\infty}$ is a diverging sequence. Moreover, using Proposition 4.3.3 we observe that
$\beta(\eta\gamma_{+},\gamma_{-},\eta\gamma_{-},\gamma_{+})$ $=\lim_{n\to\infty}[\mathtt{M}_{\rho,u}\left(\eta\gamma^{n}\eta^{-1}\gamma^{-n}\right)-\mathtt{M}_{\rho,u}\left(\eta\gamma^{n}\eta^{-1}\right)-\mathtt{M}_{\rho,u}\left(\gamma^{-n}\right)]$
$=\lim_{n\to\infty}\mathtt{M}_{\rho,u}\left(\eta\gamma^{n}\eta^{-1}\gamma^{-n}\right).$
Finally, we conclude our result using Theorem 5.1.2. ∎
Theorem 5.3.10.
Suppose $(\rho,u):\Gamma\to\mathsf{SL}_{n}\ltimes\mathfrak{sl}_{n}$ is an injective homomorphism such that $\rho$ is a Fuchsian representation and $u\in\mathsf{T}_{odd}\oplus\mathsf{T}_{2m}$ for some $1\leq m\leq\lfloor\frac{n}{2}\rfloor$. Then $(\rho,u)(\Gamma)$ does not act properly on $\mathfrak{sl}_{n}$.
Proof.
Suppose $u=u_{odd}+u_{2m}$ with $u_{odd}\in\mathsf{T}_{odd}$ and $u_{2m}\in\mathsf{T}_{2m}$.
On one hand, as $(\rho,u_{2m})(\Gamma)$ does not act properly on $\mathfrak{sl}_{n}$, using Theorem 5.1.2 we obtain a diverging sequence $\{\gamma_{n}\}_{n=0}^{\infty}$ such that $\{\mathtt{M}_{\rho,u_{2m}}(\gamma_{n})\}_{n=0}^{\infty}$ stays bounded. Suppose the attracting fixed points of $\gamma_{n}$ on $\partial_{\infty}\Gamma$ converge to $a\in\partial_{\infty}\Gamma$ and the repelling fixed points of $\gamma_{n}$ on $\partial_{\infty}\Gamma$ converge to $b\in\partial_{\infty}\Gamma$. Clearly, $a\neq b$. We choose an element $\eta\in\Gamma$ whose axis is transverse to the axis between $a$ and $b$. Hence, $\{\eta\gamma_{n}\eta^{-1}\gamma_{n}^{-1}\}_{n=0}^{\infty}$ is a diverging sequence and by Proposition 9.3 of Smilga [Smi22] we have
$\lim_{n\to\infty}\|\mathtt{M}_{\rho,u_{2m}}\left(\eta\gamma_{n}\eta^{-1}\gamma_{n}^{-1}\right)-\mathtt{M}_{\rho,u_{2m}}\left(\eta\gamma_{n}\eta^{-1}\right)-\mathtt{M}_{\rho,u_{2m}}\left(\gamma_{n}^{-1}\right)\|<\infty.$
We observe that for all $\gamma\in\Gamma$,
$\mathtt{M}_{\rho,u_{2m}}(\gamma^{-1})=\mathtt{M}_{\rho,u_{2m}}(\gamma)=\mathtt{M}_{\rho,u_{2m}}(\eta\gamma\eta^{-1}),$
and the sequence $\{\mathtt{M}_{\rho,u_{2m}}(\gamma_{n})\}_{n=0}^{\infty}$ stays bounded. Therefore, we deduce that $\{\mathtt{M}_{\rho,u_{2m}}\left(\eta\gamma_{n}\eta^{-1}\gamma_{n}^{-1}\right)\}_{n=0}^{\infty}$ also stays bounded.
On the other hand, as $\{\eta\gamma_{n}\eta^{-1}\gamma_{n}^{-1}\}_{n=0}^{\infty}$ is a diverging sequence, we again use Proposition 9.3 of Smilga [Smi22] to obtain that
$\lim_{n\to\infty}\|\mathtt{M}_{\rho,u_{odd}}\left(\eta\gamma_{n}\eta^{-1}\gamma_{n}^{-1}\right)-\mathtt{M}_{\rho,u_{odd}}\left(\eta\gamma_{n}\eta^{-1}\right)-\mathtt{M}_{\rho,u_{odd}}\left(\gamma_{n}^{-1}\right)\|<\infty.$
This time, we observe that for all $\gamma\in\Gamma$,
$\mathtt{M}_{\rho,u_{odd}}(\eta\gamma\eta^{-1})+\mathtt{M}_{\rho,u_{odd}}(\gamma^{-1})=0.$
Therefore, it follows that $\{\mathtt{M}_{\rho,u_{odd}}\left(\eta\gamma_{n}\eta^{-1}\gamma_{n}^{-1}\right)\}_{n=0}^{\infty}$ stays bounded too.
Finally, as $\mathtt{M}_{\rho,u}(\gamma)=\mathtt{M}_{\rho,u_{odd}}(\gamma)+\mathtt{M}_{\rho,u_{2m}}(\gamma)$ for all $\gamma\in\Gamma$, we conclude that the sequence $\{\mathtt{M}_{\rho,u}\left(\eta\gamma_{n}\eta^{-1}\gamma_{n}^{-1}\right)\}_{n=0}^{\infty}$ stays bounded and by Theorem 5.1.2 our result follows. ∎